# Solving Truly Massive Budgeted Monotonic POMDPs with Oracle-Guided Meta-Reinforcement Learning

### v2

@@ 3.1 Problem Statement
The primary objective of this paper is to determine a policy $\pi^{*}$ for this multi-component monotonic POMDP over a horizon $H$, that maximizes the sum of expectations of the individual times before reaching the absorbing state for each component, while adhering to the total budget $B$. We denote this maximal time $k$ by $T_{max}=\sum_{i=1}^{n}T^{i}_{max}$, where $T^{i}_{max}$ denotes the corresponding maximal time for component ii. Mathematically, the problem can be formulated as:
@@ 4.1 Budget–Concavity of the Value Function
  • $S$ is a finite state space and AA is a finite action space, as defined in Section 3.
@@
If $f(x)$ is concave on an interval II, then f(xl)f(x-l) is concave on the interval $\{y\mid y=x+l,x\in I\}$ for any constant $l$.
@@ 4.1.1 Relating $\mathbb{E}[T_{\max}]$ to the Value Function
Under reward scheme (4) each non-absorbing step contributes exactly 11 to the return; steps in state 0 contribute 0. Hence, for any budget-feasible policy $\pi$,
@@ 4.4 Optimal Policy for Multi-Component Monotonic POMDPs
We now integrate the approaches described in Section 4.2 and Section 4.3 to compute the optimal policy for an $n$-component POMDP, where $n$ is substantially large. Utilizing the random forest regressor, we efficiently approximate $\mathbb{E}[T_{\max}]$ for each component ii. Additionally, we meta-train the oracle-guided PPO agent by continuously updating the policy network’s parameters over a randomly selected subset of components and budget values. This approach allows the agent to generalize across components. This meta-trained agent is then utilized to derive the optimal policy $\pi^{i^{*}}$ for each component $i$, following the optimal budget allocation obtained from (6). Consequently, the overall policy for the multi-component POMDP is:
@@ 5.1.2 Analysis of Budget Allocation
Next, we demonstrate the effectiveness of our random forest-based budget allocation strategy. We compare it with a baseline approach that allocates budgets proportional to the ratio of a component’s replacement cost to its $\mathbb{E}[T]$. For a component $i$, we model its $\mathbb{E}[T^{i}_{\max}]$ using $\tilde{T}^{i}_{\max}$ as given in (5) (see Appendix A for justification of this exponential form). The parameters $\alpha^{i}$ and $\gamma^{i}$ can be estimated directly by considering the boundary conditions: $\gamma^{i}$ is estimated by substituting $b^{i}=0$, representing the scenario where no budget is available, and $\alpha^{i}$ is determined by substituting $b^{i}=\infty$, corresponding to the scenario of unlimited budget, where the supremum of $T^{i}_{\max}$ ($\sup_{b^{i}}T^{i}_{\max}=H=100$) is reached. We then train a random forest regressor to estimate parameter $\beta^{i}$. The training dataset is created by performing non-linear least squares regression on 11 distinct ($T^{i}_{max},b^{i}$) pairs each for 800 components. These pairs correspond to the run-averaged $T^{i}_{\max}$ values and the respective budget values $b^{i}$ from Section 5.1.1. The input to the random forest model is a vector consisting of the shape and scale factors of the Weibull distribution, which represent $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, along with the replacement and inspection costs for a given component ii. If a different distribution was used to model the transition probability, we would similarly extract the parameters, $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, for inclusion in the input vector. Figure 4 shows the prediction performance of the random forest model for a test dataset of 200 components which were not encountered during training. We see that most points on the plot are very close to the perfect prediction line and bad predictions are few in number (29 out of 200 for error threshold of $10^{-4}$). The random forest model achieves a mean squared error (MSE) = $1.8\times 10^{-8}$ for this test dataset. Note that the non-linear least squares regressor constrains $\beta^{i}$ to be $\leq 0$ and hence for some components we observe that $\beta^{i}=0$. We use this trained random forest model to estimate $\tilde{T}^{i}_{max}$ for all 1000 components. Finally, using these approximated expressions, we solve the constrained maximization problem described in (6) to obtain the appropriate budget allocation for the components.
@@ 5.1.3 Analysis of Time Complexity
Finally, we analyze the time complexity of our proposed approach for varying number of components NN. As mentioned earlier, our method comprises of four major steps:
@@
    Random Forest regression for estimating $\tilde{T}^{i}_{max}$ for each component ii.
@@ 5.2.1 Analysis of Recapitalization Policy
    Baseline (Heuristic): fixed inspect every 55 trading days; if the observed $s_{t}<20$, take recapitalize; otherwise defer.
@@
Under the shared budget $B_{\text{tot}}=15{,}000$ across $471$ components and a $120$-day evaluation horizon, we observe a consistent ordering (see Figure 7): $\textbf{Oracle}>\textbf{Oracle-guided meta-PPO}>\textbf{Baseline}>\textbf{Vanilla meta-PPO}$. The Vanilla meta-PPO tends to conserve budget and rarely recapitalizes, yielding the lowest survival time. The Baseline performs periodic inspections (every 55 trading days) and recapitalizes below the threshold but spends budget indiscriminately and misses urgent cases. By contrast, the Oracle-guided meta-PPO learns when to inspect versus defer and when to act, allocating budget to higher-value opportunities; it reliably outperforms the Baseline and closes a substantial portion of the gap to the Oracle upper bound.
@@ 5.2.2 Generalizability and Window Robustness of the Oracle-guided Meta-PPO
We vary the number of components $N\in\{5,10,20,100,471\}$. For each NN, the policy is trained on rolling $120$-day train windows and evaluated both in-sample (train) and on a held-out test window. We report average survival time (days) over r=5r{=}5 seeds; error bars denote $\pm 1$ standard deviation across seeds.
@@
(1) Both train and test curves decrease as NN grows, reflecting budget dilution and increased problem complexity. (2) Train performance is consistently above test with a modest generalization gap that tends to widen at larger NN. (3) Variability is non-negligible and generally larger at higher NN.
@@
As can be seen from Figure 8, the oracle-guided meta-PPO exhibits window robustness: trends are consistent across train windows, and the train-to-test drop remains moderate. At small NN, the effective exploration/interaction budget is limited, which can hinder learning; as NN increases, richer allocation opportunities make better use of the oracle guidance even though absolute survival time declines under a fixed total budget.
@@ Appendix A Function Approximation of $\mathbb{E}[T_{\max}]$
We model $\mathbb{E}[T^{i}_{\max}]$ as an exponential function of the budget allocated to component ii. The choice of an exponential function is motivated by its ability to capture the saturation in $\mathbb{E}[T^{i}_{\max}]$ values at higher budget levels, a result of the finite planning horizon $H$. Additionally, the exponential model accounts for non-zero $\mathbb{E}[T^{i}_{\max}]$ even when the budget is zero.

### v3

@@ 3.1 Problem Statement
The primary objective of this paper is to determine a policy $\pi^{*}$ for this multi-component monotonic POMDP over a horizon $H$, that maximizes the sum of expectations of the individual times before reaching the absorbing state for each component, while adhering to the total budget $B$. We denote this maximal time $k$ by $T_{max}=\sum_{i=1}^{n}T^{i}_{max}$, where $T^{i}_{max}$ denotes the corresponding maximal time for component $i$. Mathematically, the problem can be formulated as:
@@ 4.1 Budget–Concavity of the Value Function
  • $S$ is a finite state space and $A$ is a finite action space, as defined in Section 3.
@@
If $f(x)$ is concave on an interval $I$, then f(xl)f(x-l) is concave on the interval $\{y\mid y=x+l,x\in I\}$ for any constant $l$.
@@ 4.1.1 Relating $\mathbb{E}[T_{\max}]$ to the Value Function
Under reward scheme (4) each non-absorbing step contributes exactly $1$ to the return; steps in state 0 contribute 0. Hence, for any budget-feasible policy $\pi$,
@@ 4.4 Optimal Policy for Multi-Component Monotonic POMDPs
We now integrate the approaches described in Section 4.2 and Section 4.3 to compute the optimal policy for an $n$-component POMDP, where $n$ is substantially large. Utilizing the random forest regressor, we efficiently approximate $\mathbb{E}[T_{\max}]$ for each component $i$. Additionally, we meta-train the oracle-guided PPO agent by continuously updating the policy network’s parameters over a randomly selected subset of components and budget values. This approach allows the agent to generalize across components. This meta-trained agent is then utilized to derive the optimal policy $\pi^{i^{*}}$ for each component $i$, following the optimal budget allocation obtained from (6). Consequently, the overall policy for the multi-component POMDP is:
@@ 5.1.2 Analysis of Budget Allocation
Next, we demonstrate the effectiveness of our random forest-based budget allocation strategy. We compare it with a baseline approach that allocates budgets proportional to the ratio of a component’s replacement cost to its $\mathbb{E}[T]$. For a component $i$, we model its $\mathbb{E}[T^{i}_{\max}]$ using $\tilde{T}^{i}_{\max}$ as given in (5) (see Appendix A for justification of this exponential form). The parameters $\alpha^{i}$ and $\gamma^{i}$ can be estimated directly by considering the boundary conditions: $\gamma^{i}$ is estimated by substituting $b^{i}=0$, representing the scenario where no budget is available, and $\alpha^{i}$ is determined by substituting $b^{i}=\infty$, corresponding to the scenario of unlimited budget, where the supremum of $T^{i}_{\max}$ ($\sup_{b^{i}}T^{i}_{\max}=H=100$) is reached. We then train a random forest regressor to estimate parameter $\beta^{i}$. The training dataset is created by performing non-linear least squares regression on 11 distinct ($T^{i}_{max},b^{i}$) pairs each for 800 components. These pairs correspond to the run-averaged $T^{i}_{\max}$ values and the respective budget values $b^{i}$ from Section 5.1.1. The input to the random forest model is a vector consisting of the shape and scale factors of the Weibull distribution, which represent $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, along with the replacement and inspection costs for a given component $i$. If a different distribution was used to model the transition probability, we would similarly extract the parameters, $\mathbb{E}[T]$ and $\sigma^{2}_{\mathbb{E}[T]}$, for inclusion in the input vector. Figure 4 shows the prediction performance of the random forest model for a test dataset of 200 components which were not encountered during training. We see that most points on the plot are very close to the perfect prediction line and bad predictions are few in number (29 out of 200 for error threshold of $10^{-4}$). The random forest model achieves a mean squared error (MSE) = $1.8\times 10^{-8}$ for this test dataset. Note that the non-linear least squares regressor constrains $\beta^{i}$ to be $\leq 0$ and hence for some components we observe that $\beta^{i}=0$. We use this trained random forest model to estimate $\tilde{T}^{i}_{max}$ for all 1000 components. Finally, using these approximated expressions, we solve the constrained maximization problem described in (6) to obtain the appropriate budget allocation for the components.
@@ 5.1.3 Analysis of Time Complexity
Finally, we analyze the time complexity of our proposed approach for varying number of components $N$. As mentioned earlier, our method comprises of four major steps:
@@
    Random Forest regression for estimating $\tilde{T}^{i}_{max}$ for each component $i$.
@@ 5.2.1 Analysis of Recapitalization Policy
    Baseline (Heuristic): fixed inspect every $5$ trading days; if the observed $s_{t}<20$, take recapitalize; otherwise defer.
@@
Under the shared budget $B_{\text{tot}}=15{,}000$ across $471$ components and a $120$-day evaluation horizon, we observe a consistent ordering (see Figure 7): $\textbf{Oracle}>\textbf{Oracle-guided meta-PPO}>\textbf{Baseline}>\textbf{Vanilla meta-PPO}$. The Vanilla meta-PPO tends to conserve budget and rarely recapitalizes, yielding the lowest survival time. The Baseline performs periodic inspections (every $5$ trading days) and recapitalizes below the threshold but spends budget indiscriminately and misses urgent cases. By contrast, the Oracle-guided meta-PPO learns when to inspect versus defer and when to act, allocating budget to higher-value opportunities; it reliably outperforms the Baseline and closes a substantial portion of the gap to the Oracle upper bound.
@@ 5.2.2 Generalizability and Window Robustness of the Oracle-guided Meta-PPO
We vary the number of components $N\in\{5,10,20,100,471\}$. For each $N$, the policy is trained on rolling $120$-day train windows and evaluated both in-sample (train) and on a held-out test window. We report average survival time (days) over r=5r{=}5 seeds; error bars denote $\pm 1$ standard deviation across seeds.
@@
(1) Both train and test curves decrease as $N$ grows, reflecting budget dilution and increased problem complexity. (2) Train performance is consistently above test with a modest generalization gap that tends to widen at larger $N$. (3) Variability is non-negligible and generally larger at higher $N$.
@@
As can be seen from Figure 8, the oracle-guided meta-PPO exhibits window robustness: trends are consistent across train windows, and the train-to-test drop remains moderate. At small $N$, the effective exploration/interaction budget is limited, which can hinder learning; as $N$ increases, richer allocation opportunities make better use of the oracle guidance even though absolute survival time declines under a fixed total budget.
@@ Appendix A Function Approximation of $\mathbb{E}[T_{\max}]$
We model $\mathbb{E}[T^{i}_{\max}]$ as an exponential function of the budget allocated to component $i$. The choice of an exponential function is motivated by its ability to capture the saturation in $\mathbb{E}[T^{i}_{\max}]$ values at higher budget levels, a result of the finite planning horizon $H$. Additionally, the exponential model accounts for non-zero $\mathbb{E}[T^{i}_{\max}]$ even when the budget is zero.